a force of 2 n acts for 5 seconds on a particla of mass 0.5 kg initially at rest calculate the distance moved by the particle in this 5 seconds and next 5 seconds

Grade:9

1 Answers

Aarti Gupta
9 years ago
Given – Force,F = 2N
Time,t = 5s
Mass of particle,m = O.5 Kg
Initial velocity,u = 0 m/s
Distance moved in this 5s, S = ?
Solution -
Applying the equation,
S = ut + ½ at2
Now F= m*a ,so a = F/m = 2/0.5 = 4m/s2
Sustituting the values,
S = 0*(5)2 + ½ *4* (5)2
S = 50m
For next 5s total time = 10s
Thus, S = ½ *4 * (10)2
S = 200m